# Lossy Compression with Data, Perception, and Classification Constraints

### v1

@@ Appendix C Proof of Theorem 3
Recall the following information rate-perception classification (RPC) function:
$R(P,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$ (3)
s.t. $d(p_{X},p_{\hat{X}})\leq P,$ (3a)
$H(S|\hat{X})\leq C,$ (3b)
where $S$ is the only classification variable.
Consider a Bernoulli source $X$ and a classification variable $S$ with binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$). Similar to the RDC case, if $C<H(S_{1})$, by data-processing inequality[15, Theorem 2.8.1], the problem is infeasible. In the following discussion, we assume $C\geq H(S_{1})$.
To derive the optimal rate, we need to find the optimal solution $p_{a}=P(\hat{X}=0|X=0)$ and $p_{b}=P(\hat{X}=0|X=1)$ to binary RPC problem (3). From the BSC assumption, we can obtain that the marginal distribution of $X$ is $\text{Pr}(X=1)=\frac{a-p_{1}}{1-2p_{1}}\triangleq b$.
@@
$d_{TV}(p_{X},p_{\hat{X}})$ $=\frac{1}{2}\sum_{z\in\{0,1\}}|p_{X}(z)-p_{\hat{X}}(z)|$
$=|-bp_{b}+(1-b)(1-p_{a})|.$
Now we express $H(S|\hat{X})$ as a function of $p_{a},p_{b}$. By the definition of conditional entropy and $S-X-\hat{X}$, we have
$H(S|$ $\hat{X})\!\!=\!-(bp_{1}p_{b}\!\!+\!\!(\!1\!-\!b\!)\!(\!1\!-\!p_{1}\!)p_{a})\!\log\!\Big{(}\!\frac{bp_{1}p_{b}\!\!+\!\!(\!1\!\!-\!\!b)\!(1\!\!-\!p_{1}\!)p_{a}}{bp_{b}\!\!+\!\!(1\!\!-\!b)p_{a}}\!\Big{)}$
-- $\!(bp_{1}\!(\!1\!-\!p_{b}\!)\!\!+\!\!(1\!-\!b)(1\!-\!p_{1})(1\!-\!p_{a}\!))$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \cdot\log\Big{(}\!\frac{bp_{1}\!(1\!-\!p_{b})\!\!+\!(1\!\!-\!b)(1\!\!-\!p_{1})(\!1\!\!-\!p_{a})}{b(\!1\!-\!p_{b})\!+\!(1\!-\!b)p_{a}}\Big{)}$
-- $\!(b(1\!\!-\!p_{1})p_{b}\!+\!(1\!\!-\!b)p_{1}p_{a})\cdot\!\log\Big{(}\!\frac{b(1\!\!-\!p_{1}\!)p_{b}\!+\!(1\!-\!b)p_{1}p_{a}}{bp_{b}\!+\!(1\!-\!b)p_{a}}\!\Big{)}$
-- $(b(1\!-\!p_{1})(1\!\!-\!p_{b})\!+\!(1\!\!-\!b)p_{1}(1\!\!-\!p_{a}))$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \cdot\log\Big{(}\!\frac{b(1\!-\!p_{1})(1\!-\!p_{b})\!+\!(1\!-\!b)p_{1}(1\!-\!p_{a})}{b(1\!-\!p_{b})\!+\!(1\!-\!b)p_{a}}\!\Big{)}$
$\triangleq f(p_{a},p_{b}).$
@@
$p_{b}=\frac{(1-b)(1-p_{a})}{b},$ (10)
$p_{a}=g^{-1}(C),$ (11)
where $g^{-1}:[0,1]\rightarrow[\frac{1}{2},1]$ denote the inverse function of
$g(p_{a})\triangleq f(p_{a},\frac{(1-b)(1-p_{a})}{b}).$ (12)
@@
$\frac{\text{d}g(p_{a})}{\text{d}p_{a}}$ $=-(\log A+1)\frac{\partial A}{\partial p_{a}}-(\log B+1)\frac{\partial B}{\partial p_{a}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ -(\log C+1)\frac{\partial C}{\partial p_{a}}-(\log D+1)\frac{\partial D}{\partial p_{a}}$
$=-(1\!-\!2p_{1})(1\!-\!b)\log A\!-\!(2p_{1}\!-\!1)(1\!-\!b)\log B$
$\ \ \ \ \ \ \ \!-\!(2p_{1}\!-\!1)(1\!-\!b)\log C\!-\!(1\!-\!2p_{1})(1\!-\!b)\log D$
$=(1-2p_{1})(1-b)\log\frac{B\cdot C}{A\cdot D}.$
@@
Then we can see that the pair of $(p_{a},p_{b})$ defined in (11) and (10) makes the perception constraint (3a) always inactive, since for $p_{b}=\frac{(1-b)(1-p_{a})}{b}$, we have
@@
At the same time, $H(S|\hat{X})=f(p_{a},\frac{(1-b)(1-p_{a})}{b})=g(p_{a})=C$ for the choice of $p_{a}=g^{-1}(C)$. By Mrs. Gerber’s Lemma[16], when $H(S|\hat{X})\leq C$, we have $H(X|\hat{X})\leq H(C_{1})$, with $C_{1}$ defined as $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$. The equality $H(X|\hat{X})=H(C_{1})$ holds when $H(S|\hat{X})=C$. Hence, such choice of $(p_{a},p_{b})$ achieves the lower bound of $I(X;\hat{X})=H(X)-H(C_{1})$ with the perception constraint always inactive.
@@ Appendix D Proof of Theorem 4
Consider the RPC problem as below,
$R(P,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$
s.t. $d_{\text{KL}}(p_{X},p_{\hat{X}})\leq P,$
$H(S|\hat{X})\leq C,$
where $(X,S)$ are jointly Gaussian variable with covariance $\text{Cov}(X,S)=\theta_{1}$. Similar to [9, 17], the optimal reconstruction $\hat{X}$ should also be jointly Gaussian with $X$, thus the parameters to be optimized are expectation $\mu_{\hat{x}}$, variance $\sigma_{\hat{x}}^{2}$, and covariance $\text{Cov}(X,\hat{X})=\theta_{2}$.
First we compute the KL-divergence of two Gaussian varibales $X\sim\mathcal{N}(\mu_{x},\sigma^{2}_{x})$ and $\hat{X}\sim\mathcal{N}(\mu_{\hat{x}},\sigma^{2}_{\hat{x}})$. By definition,
$d_{\text{KL}}(p_{X},p_{\hat{X}})$ $=\int p_{X}(x)\log\frac{p_{X}(x)}{p_{\hat{X}}(x)}dx$
[…]
$=\frac{1}{2}\log(2\pi\sigma^{2}_{\hat{x}})+\frac{1}{\sigma_{\hat{x}}^{2}}(\sigma^{2}_{x}+\mu^{2}_{x})+\frac{\mu_{\hat{x}}}{\sigma^{2}_{\hat{x}}}\mu_{x}+\frac{\mu^{2}_{\hat{x}}}{2\sigma^{2}_{\hat{x}}}$
$=\frac{1}{2}\log(2\pi\sigma^{2}_{\hat{x}})+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}}{2\sigma_{\hat{x}}^{2}}.$
Thus,
$d_{\text{KL}}(p_{X},p_{\hat{X}})$ $=\!-\frac{1}{2}\!\log(2\pi e\sigma_{x}^{2})\!+\!\frac{1}{2}\log(2\pi\sigma^{2}_{\hat{x}})\!+\!\frac{(\mu_{x}\!-\!\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}\!+\!\frac{\sigma_{x}^{2}}{2\sigma_{\hat{x}}^{2}}$
$=\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}},$
where we take the natural base of logarithm.
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[15, Chapter 8], as well as the KL divergence of two Gaussian varibales, the RPC problem can be represented by
@@
First, we can observe that we can set $\mu_{\hat{x}}=\mu_{x}$ without any effect of the objective function and the constraint (13b), while reducing the perception constraint (13a).
Similar to the RDC case, by (13b) we have
$\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}\geq\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C}),$
which gives us a lower bound of the objective function $I(X;\hat{X})$:
$-\frac{1}{2}\log(1\!-\!\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ $\geq\!-\frac{1}{2}\log\Big{(}1\!-\!\frac{1}{\sigma_{x}^{2}}\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1\!-\!e^{-2h(S)+2C})\Big{)}$
$=-\frac{1}{2}\log\Big{(}1-\frac{\sigma_{s}^{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\Big{)},$
since $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ is an increasing function of $\frac{\theta_{2}^{2}}{\sigma^{2}_{\hat{x}}}$. The lower bound is achieved when $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$.
Now we show that the lower bound could be achieved with the constraint (13a) always inactive.
@@
$\sigma_{\hat{x}}=\sigma_{x},$ (14)
$\theta_{2}^{2}=\frac{\sigma_{s}^{2}\sigma_{x}^{6}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C}).$ (15)
@@
Now we verify that the solution (15) is valid. Since the correlation $\rho=\frac{\theta_{2}}{\sigma_{x}\sigma_{\hat{x}}}$ should have value between -1 and 1, we require that
$\rho^{2}=\frac{\theta_{2}^{2}}{\sigma^{2}_{x}\sigma^{2}_{\hat{x}}}=\frac{\sigma_{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\leq 1,$
which gives us
$C\geq\frac{1}{2}\log\Big{(}1-\frac{\theta_{2}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}}\Big{)}+h(S).$

### v2

@@ Appendix C Proof of Theorem 3
Recall the RPC function in (3) for a Bernoulli source $X$ and a classification variable $S$ with binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$). Similar to the RDC case, if $C<H(S_{1})$, by data-processing inequality[17, Theorem 2.8.1], the problem is infeasible. In the following discussion, we assume $C\geq H(S_{1})$.
Converse: By Mrs. Gerber’s Lemma proved in [18], when $H(S|\hat{X})\leq C$, we have $H(X|\hat{X})\leq H(C_{1})$, where $C_{1}\triangleq\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$. Note that $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}\leq\frac{\frac{1}{2}-p_{1}}{1-2p_{1}}=\frac{1}{2}$. Therefore, we have $I(X;\hat{X})=H(X)-H(X|\hat{X})\geq H(b)-H(C_{1})$.
Achievability: To achieve the optimal rate, we need to find the optimal solution $p_{a}=P(\hat{X}=0|X=0)$ and $p_{b}=P(\hat{X}=0|X=1)$. From the BSC assumption, we can obtained the marginal distribution of $X$ is $\text{Pr}(X=1)=\frac{a-p_{1}}{1-2p_{1}}\triangleq b$.
@@
$d_{TV}(p_{X},p_{\hat{X}})$ $=|-bp_{b}+(1-b)(1-p_{a})|.$
By the definition of conditional entropy and $S-X-\hat{X}$, we also express $H(S|\hat{X})$ as a function of $p_{a},p_{b}$, and denote $f(p_{a},p_{b})\triangleq H(S|\hat{X})$.
@@
$p_{b}=\frac{(1-b)(1-p_{a})}{b},~{}p_{a}=g^{-1}(C),$ (9)
where $g^{-1}:[0,1]\rightarrow[\frac{1}{2},1]$ denote the inverse function of $g(p_{a})\triangleq f(p_{a},(1-b)(1-p_{a})/b)$.
@@
$\frac{\text{d}g(p_{a})}{\text{d}p_{a}}$ $=(1-2p_{1})(1-b)\log\frac{B\cdot C}{A\cdot D}.$
@@
Then we can see that the pair of $(p_{a},p_{b})$ defined in (9) makes the perception constraint (3a) always inactive, since for $p_{b}=\frac{(1-b)(1-p_{a})}{b}$, we have
@@
At the same time, $H(S|\hat{X})=f(p_{a},\frac{(1-b)(1-p_{a})}{b})=g(p_{a})=C$ for the choice of $p_{a}=g^{-1}(C)$.
By Mrs. Gerber’s Lemma[18], when $H(S|\hat{X})\leq C$, we have $H(X|\hat{X})\leq H(C_{1})$, with $C_{1}$ defined as $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$. The equality $H(X|\hat{X})=H(C_{1})$ holds when $H(S|\hat{X})=C$. Hence, such choice of $(p_{a},p_{b})$ achieves the lower bound of $I(X;\hat{X})=H(X)-H(C_{1})$ with the perception constraint always inactive.
@@ Appendix D Proof of Theorem 4
Consider the RPC function in (3) where $(X,S)$ are jointly Gaussian variable with covariance $\text{Cov}(X,S)=\theta_{1}$. Similar to [11, 19], the optimal reconstruction $\hat{X}$ should also be jointly Gaussian with $X$, thus the parameters to be optimized are expectation $\mu_{\hat{x}}$, variance $\sigma_{\hat{x}}^{2}$, and covariance $\text{Cov}(X,\hat{X})=\theta_{2}$.
The KL-divergence of two Gaussian varibales $X\sim\mathcal{N}(\mu_{x},\sigma^{2}_{x})$ and $\hat{X}\sim\mathcal{N}(\mu_{\hat{x}},\sigma^{2}_{\hat{x}})$ is $d_{\text{KL}}(p_{X},p_{\hat{X}})=\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}},$ where we take the natural base of logarithm.
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[17, Chapter 8], as well as the KL divergence of two Gaussian varibales, the RPC problem can be represented by
@@
First, it is observed that we can set $\mu_{\hat{x}}=\mu_{x}$ without any effect of the objective function and the constraint (10b), while reducing the perception constraint (10a).
Converse: Similar to the RDC case, by (10b) we have $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}\geq\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$, which gives us a lower bound of the objective function $I(X;\hat{X})-\frac{1}{2}\log\Big{(}1-\frac{\sigma_{s}^{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\Big{)}$, since $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ is an increasing function of $\frac{\theta_{2}^{2}}{\sigma^{2}_{\hat{x}}}$. The lower bound is achieved when $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$.
Achievability: Now we show that the lower bound could be achieved with the constraint (10a) always inactive.
@@
$\sigma_{\hat{x}}=\sigma_{x},~{}\theta_{2}^{2}=\frac{\sigma_{s}^{2}\sigma_{x}^{6}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C}).$ (11)
@@
Now we verify that the solution (11) is valid. Since the correlation $\rho=\frac{\theta_{2}}{\sigma_{x}\sigma_{\hat{x}}}$ should have value between -1 and 1, we require that $\rho^{2}=\frac{\theta_{2}^{2}}{\sigma^{2}_{x}\sigma^{2}_{\hat{x}}}=\frac{\sigma_{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\leq 1,$ which gives us $C\geq\frac{1}{2}\log\Big{(}1-\frac{\theta_{2}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}}\Big{)}+h(S)$.